put the equation in slope intercept form, determine the slope and y-intercept and graph equation. 2y-x=8. m= and B=
Answer by mananth(16949)

2y-x=8
2y=x+8
y=x/2+4
m=slope = 1/2
b= 4
when x= 0,y=4 (0,4)
when y =0,x=-8 (-8,0)